OpenStudy (nathalyn):

8 = 3(3x+8) - x Please help me !

pooja195 (pooja195):

8=3(3x+8)−x 8=(3)(3x)+(3)(8)+−x(Distribute) 8=9x+24+−x 8=(9x+−x)+(24)(Combine Like Terms) 8=8x+24 8=8x+24

pooja195 (pooja195):

Flip the equation. 8x+24=8 Subtract 24 from both sides. 8x+24−24=8−24 8x=−16 divide 8 from both sides 8/8=1 so it cancels out and we are left with x -16/8=-2 x=-2

OpenStudy (nathalyn):

WOW THANNK YOU!

OpenStudy (nathalyn):

I wish i was intelligent! :(

pooja195 (pooja195):

Everyone is intelligent in their own way :D

OpenStudy (nathalyn):

:)

OpenStudy (mrnood):

maybe you would become more confident in your own intelligence and capability if you tried to do some of the work yourself. @pooja195 you really don't help by working out the complete answer like that it is my bet that faced with an almost identical question the poster would not be closer to doing it now than before this exercise

pooja195 (pooja195):

@MrNood i know how to teach people. ive been here long enough to know.

pooja195 (pooja195):

If i wanted i could have just given the answer

OpenStudy (nathalyn):

i go with @pooja195 :)

pooja195 (pooja195):

But i didnt i showed my work so the asker understands and can go back and look at the steps

OpenStudy (mrnood):

but it is better to tell them the steps and ask THEM to actually do it step bu step for example you could have said " the first step is to distribute the 3 into the brackets" all your poster probably took from your long pot was "x=-2" "I DO and I understand" is the way to learn.- NOT a long process of someone else proving that THEY know how to do it.

OpenStudy (nathalyn):

okay let me see what your teach

OpenStudy (mrnood):

well - 2 things - first feel free to look at my profile and the many hundreds of questions I have GUIDED people to the answer - with THEIR OWN effort. Second - if you think that was good teaching can you do this? 13-x= 6(4x-2) it is EXACTLY the same principle as above, but requires you to have understood the steps taken to get the answer

OpenStudy (mrnood):

maybe you'd show your steps in the same way that pooja did above